OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. A. 84 B. 7 C. -12

OpenStudy (the_fizicx99):

y = kx is the direct variation formula -7 = -1k /-1 /-1 k = 7 So -84 =7k /7 /7 k = -12 -84 = 7(-12) -84 = -84 <-- true